stopscammingpeople:

-(-(-)(-)(-10x))=-5 solve for x.

snowflake0531:

two negative signs can be crossed out at once \(-(-(-)(-)(-10x))=-5\\-(-(-10x))=-5\\-10x=-5\\10x=5\\\frac{10}{10}x=\frac{5}{10}\\x=\frac{1}{2}\)

snowflake0531:

two negative signs can be crossed out at once because they make a positive \[-(-(-(-(-10x))=-5\\-(-(-10x))=-5\\-10x=-5\\10x=5\\\frac{10}{10}x=\frac{5}{10}\\x=\frac{1}{2}\]